### v1

# AnyCharV: Bootstrap Controllable Character Video Generation with Fine-to-Coarse Guidance

Abstract
Character video generation is a significant real-world application focused on producing high-quality videos featuring specific characters. Recent advancements have introduced various control signals to animate static characters, successfully enhancing control over the generation process. However, these methods often lack flexibility, limiting their applicability and making it challenging for users to synthesize a source character into a desired target scene. To address this issue, we propose a novel framework, AnyCharV, that flexibly generates character videos using arbitrary source characters and target scenes, guided by pose information. Our approach involves a two-stage training process. In the first stage, we develop a base model capable of integrating the source character with the target scene using pose guidance. The second stage further bootstraps controllable generation through a self-boosting mechanism, where we use the generated video in the first stage and replace the fine mask with the coarse one, enabling training outcomes with better preservation of character details. Experimental results demonstrate the effectiveness and robustness of our proposed method. Our project page is https://anycharv.github.io.
1 Introduction
Recent advancements in diffusion modeling have significantly advanced character video generation for various applications (Blattmann et al., 2023; Ho et al., 2022; Hong et al., 2022; Wu et al., 2023b; Zhang et al., 2023b; Wang et al., 2024f; Wu et al., 2024). While these approaches achieve remarkable success with guidance from text and image priors, they often fail to preserve the intrinsic characteristics of the character, such as nuanced appearance and complex motion. Additionally, these methods lack the capability for users to exert individual control during the character video generation process, making it difficult to modify the character’s motion and the surrounding scene. These limitations hinder the practical application and development of character video generation technologies.
Several recent works aim to improve the controllability of character video generation by incorporating additional conditions, such as pose (Hu, 2024; Xu et al., 2024; Zhang et al., 2024; Zhu et al., 2025), depth (Xing et al., 2025; Zhu et al., 2025), and normal maps (Zhu et al., 2025). However, these methods primarily focus on animating characters within fixed background scenes or customizing characters against random backgrounds, lacking the capability for precise background control and often introducing artifacts. These limitations reduce the flexibility of character video generation and imposes significant constraints on practical applications. Directly replacing a character within a given scene is crucial for applications such as art creation and movie production. Some related works attempt to address this issue using 3D priors (Men et al., 2024; Viggle, 2024) and naive 2D representations (Zhao et al., 2023; Qiu et al., 2024). However, characters constructed from 3D information often lack realistic interaction appearances, while 2D prior-driven characters suffer from temporal inconsistencies and noticeable jittering.
In this work, we aim to develop a flexible and high-fidelity framework for controllable character video generation. Given the complexity and challenges of this task, we propose a novel two-stage approach with fine-to-coarse guidance, AnyCharV, consisting of a self-supervised composition stage and a self-boosting training stage. In the first stage, we construct a base model to integrate the reference character and the target scene, using fine guidance from the target character’s segmentation mask and the pose as a conditional signal. During this phase, the model learns to accurately compose the source character and target scene with precise spatial and temporal control, akin to existing methods that guide the model with controlled elements. However, due to shape differences between the source and target characters, the generated video often appears blurry, contains artifacts, and lacks realism. To address this critical issue, which is a significant drawback of current research, we introduce a self-boosting training stage. In this stage, we generate multiple source-target video pairs using the model from the first stage and employ these synthesized videos for bootstrap training. These video pairs share the same background scene, with the source and target characters in identical poses. Instead of using a fine target character segmentation mask, we utilize a bounding box mask to provide coarse guidance for the character area. By focusing solely on the source character and not separately feeding the model the target video scene, we enable better character control and detail recovery during self-boosting training. This approach ensures that the details of the source character are better preserved during subsequent inference, as shown in Figure 1. In summary, our contributions are as follows:
We propose a novel framework AnyCharV for controllable character video generation that employs fine-to-coarse guidance. This approach enables the seamless integration of any source character into a target video scene, delivering flexible and high-fidelity results.
We achieve the composition of the source character and target video scene in a self-supervised manner under the guidance of a fine segmentation mask, introducing a simple yet effective method for controllable generation.
We develop a self-boosting strategy that leverages the interaction between the target and source characters with guidance from a coarse bounding box mask, enhancing the preservation of the source character’s details during inference.
Our method demonstrates superior generation results, outperforming previous open-sourced state-of-the-art approach both qualitatively and quantitatively.
2 Related Work
2.1 Controllable Video Generation
Controlling the process of video generation is crucial and has been extensively studied. Recent research efforts in this direction often rely on introducing additional control signals, such as depth maps (Chai et al., 2023; Zhang et al., 2023c; Wang et al., 2024b), canny edges (Zhang et al., 2023c), text descriptions (Zhang et al., 2023b), sketch maps (Wang et al., 2024b), and motions (Wang et al., 2024b). These works achieve control by incorporating these signals into the video generation model, resulting in the desired outcomes. Meanwhile, other approaches focus on learning high-level feature representations related to appearance (He et al., 2024; Wang et al., 2024f; Wei et al., 2024) and motion (Wu et al., 2024; Yang et al., 2024a; Zhao et al., 2025) within video diffusion models. These methods aim to customize the video generation by learning the identity of subjects and motions. However, the controls provided by these approaches are often too coarse to meet the stringent requirements for controllable character video generation, where fine-grained control over both character and scene is essential.
2.2 Character Video Generation
Character video generation has achieved remarkable results recently with the development of diffusion models (Ho et al., 2020; Song et al., 2020; Lu et al., 2022). Generating a character video with high-fidelity is challenging due to the complex character motion and various appearance details. To achieve high-quality generation, some approaches try to train a large model upon vast amounts of data (Liu et al., 2024; Yang et al., 2024b; Li et al., 2024; Guo et al., 2023; Blattmann et al., 2023; Bao et al., 2024). In this case, the character video can be directly synthesized by the input text prompt or reference image. Nevertheless, these direct methods introduce significant randomness when generating the characters, resulting in many artifacts and unsmooth motions. To tackle these issues, recent researches try to guide the character video generation with 2D pose sequence (Hu, 2024; Xu et al., 2024; Peng et al., 2024; Chang et al., 2023, 2024; Wang et al., 2024c; Zhai et al., 2024; Wang et al., 2024a), depth maps (Zhu et al., 2025), and 3D normal (Zhu et al., 2025). However, these approaches are limited to animating the given image within its original background scene, which restricts their applicability in diverse scenarios.
Recent works have proposed more flexible settings by combining arbitrary characters with arbitrary background scenes. An early work, Make-A-Protagonist (Zhao et al., 2023), modifies the foreground character using image prompts and mask supervision based on the Stable Diffusion model (Rombach et al., 2022). However, this approach suffers from poor temporal consistency in the generated videos, and the background scene can change unexpectedly, which is undesirable for users. MIMO (Men et al., 2024) addresses this by decoupling the foreground character and background scene for further composition during video generation. Although MIMO’s approach is based on 3D representations, it can result in unrealistic video generation. Additionally, inaccuracies in conditioned 3D motion, occlusion, and structure modeling can degrade performance. Viggle (Viggle, 2024), which is also based on 3D representations, encounters similar challenges. MovieCharacter (Qiu et al., 2024) introduces a tuning-free framework for video composition using pose-guided character animation and video harmonization (Guerreiro et al., 2023). However, MovieCharacter struggles with complex motions and interactions. In contrast to these methods, we address the composition problem using a self-supervised training framework with fine-to-coarse guidance driven by 2D pose guidance, enhancing composition capabilities and improving generation quality.
3 Method
We aim to generate a video featuring our desired character, guided by a target driving video that provides character motion, background scene, and human-object interaction information. This task presents several significant challenges. Firstly, the appearance of the given character must be well-preserved to ensure that the generated character is vivid and high-fidelity. Secondly, the model should accurately animate the character to follow the motion provided by the target driving video. Thirdly, to ensure the seamless integration of the generated character into the target scene within the corresponding region, the generation model is required to learn the complex human-object interactions and vividly represent these connections. To address these challenges, we design a two-stage training framework AnyCharV with fine-to-coarse guidance. In the following, we successively describe the preliminary of video diffusion model in Section 3.1, the first stage of our framework about training a base model for controllable character video generation with fine mask guidance in Section 3.2, and the second stage of our framework on how to improve the basic generator via our proposed self-boosting training strategy with coarse mask guidance in Section 3.3. An overview of our proposed framework is illustrated in Figure 2.
3.1 Preliminary: Video Diffusion Model
Video diffusion models (VDMs) (Chen et al., 2024; Blattmann et al., 2023; Yang et al., 2024b) synthesize videos by successively refining randomly sampled Gaussian noise . This procedure parallels the reversal of a fixed-length Markov chain. Through iterative denoising, VDMs capture the temporal relationships embedded in video data. Concretely, at each timestep , a video diffusion model estimates the noise given an image condition . The training process optimizes a reconstruction objective:
where denotes the latent representation of the video, with as the batch size, the sequence length, and the spatial dimensions, and the latent dimensionality. Here, is the model-estimated noise. The noisy intermediate state is obtained by combining ground-truth features with noise as , where is a diffusion parameter.
3.2 Self-Supervised Composition with Fine Guidance
Naive Self-Supervised Strategy.
To train our model, it is impractical to collect a large number of video pairs that feature the same character motion and scene but with different characters. Therefore, we propose a naive self-supervised strategy to make our training feasible. Supposing we have a training video , we can easily sample a reference image and a target driving video from . Given , we extract the corresponding character segmentation mask by a segmentation model, such as SAM (Kirillov et al., 2023). To this end, we can train the model to generate with the reference character and the target scene , where is the complementary part of the mask , and is a all-ones matrix. During training, the target scene is concatenated with the noisy target video and fed into the denoising UNet, as shown in Figure 2.
Reference Prior Preservation.
As mentioned before, preserving the appearance and identity of the reference character is vital in our studied problem. Thus, in our design, the identity of the reference character is introduced via two pathways. Firstly, the reference image is encoded by a pre-trained CLIP (Radford et al., 2021) image encoder and conditioned on the cross-attention layers. However, such CLIP-style feature injection leads to information loss, due to low resolution () and coarse feature representation. Secondly, with the success of previous character animation works (Hu, 2024; Xu et al., 2024), we utilize a ReferenceNet to better encode the character appearance. This ReferenceNet only focus on extracting the spatial features, without temporal layers. The spatial features from ReferenceNet are concatenated with spatial features from denoising UNet in the spatial attention layers for self-attention modeling. We remove the background scene in the reference image to avoid introducing redundant context information.
Fine Mask Supervision.
With the character appearance preserved, the character should be composed with the target scene in the correct region. However, naively taking the reference character and the target scene will lead to high randomness of the composition process. To this end, we further control the generation by introducing the target character segmentation mask to guide the generation region. This fine segmentation mask is directly concatenated on the noisy target video together with the target scene as a whole. However, we find that such fine mask will introduce some harmful shape information, especially when there is large difference between the reference and target characters regarding the appearance shape. In this case, to reduce this redundant information input, we apply strong augmentation onto the input fine character segmentation mask to make it have irregular mask border. Following, the scene is obtained by masking with the augmented mask .
Pose Guidance.
Smoothly animating the reference character in the target scene is required for compositional generation. Recent works (Hu, 2024; Xu et al., 2024) has achieved great success on animating arbitrary character with arbitrary pose sequence. Similar with them, we build a lightweight pose guider to embed the target character pose , consisting of 4 convolution layers, and the embedded pose images are added to the noisy target video latents. The pose can be extracted from the target video via a pose detector, such as DWPose (Yang et al., 2023).
3.3 Self-Boosting Training with Coarse Guidance
With the self-supervised composition strategy introduced in Section 3.2, we develop a base model for integrating a reference character with a target driving video. However, the guidance from the fine segmentation mask can distort the shape of the reference character, resulting in diminished reference identity preservation and overall video quality. The proposed mask augmentation mechanism still falls short of achieving optimal results. To address this, we introduce a self-boosting strategy to mitigate the negative impact of the fine mask shape. Instead of using a precise segmentation mask of the target character for training, we rely on a coarse bounding box mask to indicate the approximate region. Additionally, there’s no need to extract the target scene separately, as the target video itself can provide the necessary context of the background scene, avoiding excessive shape borders. Then we show how to better learn this compositional generation with coarse bounding box mask and original target video, assisted by self-boosting training with the generated videos in the first stage.
Pair Construction.
As mentioned in the beginning of Section 3.2, we lack video pairs with the same character motion and scene but different characters to train our model. Given a base model developed using the self-supervised composition strategy in Section 3.2, it becomes possible to create such pairs through generation. We create the data pairs via drawing samples in the original dataset, where we choose one video for reference character and another for driving video, inspired by the generative enhancement technique (Zhu et al., 2024). Specifically, we randomly sample a frame from the first video to serve as the reference character image and regard another video as the target driving video . Given and , we can generate a result video, where we name it as reference driving video . To this end, we obtain a video pair , in which the character motion, scene, and human-object interaction are all the same while they features different characters. We randomly sample videos from our training data and build 64,000 video pairs following the above process.
Self-Boosting Training.
To better preserve the identity of the reference character and eliminate the harmful impact of the mask shape, we construct a number of video pairs showing different characters with the same motion and scene information for model training. During training, instead of separately extracting the scene from the target driving video , we take the generated reference driving video as the input to provide the scene information. To help the model localize the corresponding area of the target character, we provide a coarse bounding box mask as the coarse guidance. Given the reference driving video and coarse bounding box mask , the model learns to build interactions between the reference and target characters, as well as integrate the reference character with the background scene. By loosening the mask guidance from fine to coarse, we mitigate the negative influence of the fine mask shape incurred in Section 3.2, resulting in more natural generated videos that better preserve the details of the reference character.
3.4 Training and Inference
Training Strategy.
The training of our model is in a two-stage manner. In the first stage, we train the denoising UNet, ReferenceNet, and pose guider and fix VAE (Kingma, 2013) and CLIP image encoder (Radford et al., 2021). The training objective is as following:
In the second stage, we load the trained weights of the denoising UNet, ReferenceNet, and pose guider in the stage1 and only finetune the denoising UNet. The training loss is
where is the random noise, is the noise prediction from , is the latent of target driving video , is the 2D pose sequence extracted from reference driving video , and is the sampling timestep.
Inference.
During inference, we only need a reference image and a target driving video to generate the desired output. The target character’s pose and bounding box mask are extracted in advance. With our proposed two-stage training mechanism, our model can produce high-fidelity videos that accurately preserve the identity of the reference character and the target background scene. The generated videos exhibit smooth motions and natural human-object interactions.
4 Experiment
4.1 Experimental Setting
We build a character video dataset called CharVG to train our proposed model, which contains 5,482 videos from the Internet with various characters. These videos range from 7 to 47 seconds in length. The ReferenceNet and denoising UNet are both initialized from Stable Diffusion 1.5 (Rombach et al., 2022), where the motion module in the denoising UNet is initialized from AnimateDiff (Guo et al., 2024). The pose guider is initialized from ControlNet (Zhang et al., 2023a). We conduct both qualitative and quantitative evaluation for our method. For quantitative evaluation, we collect 10 character images and 10 target driving videos from the internet, then generate videos with every image-video pair, which results in 100 evaluation videos. We adopt FVD (Unterthiner et al., 2018), Dover++ (Wu et al., 2023a), and CLIP Image Score (Radford et al., 2021) to evaluate the generation quality. Please refer to Section A for detailed experimental settings.
| Method | FVD | DOVER++ | CLIP-I |
| w/o mask aug. | 2719.46 | 51.09 | 72.68 |
| w/o self-boosting | 2784.62 | 51.24 | 72.41 |
| AnyCharV (ours) | 2582.59 | 52.20 | 73.15 |
| Mask Type | FVD | DOVER++ | CLIP-I |
| Box & Seg. | 2971.43 | 50.91 | 72.49 |
| Box & Box | 2806.74 | 51.28 | 72.71 |
| Seg. & Seg. | 2759.62 | 51.72 | 72.60 |
| Seg. & Box | 2582.59 | 52.20 | 73.15 |
4.2 Comparison
We compare our approach with three recent state-of-the-art methods, including Make-A-Protagonist (Zhao et al., 2023), Viggle (Viggle, 2024), and MIMO (Men et al., 2024).
Qualitative Results.
As shown in Figure 3, Make-a-Protagonist performs worst as for lacking effective spatial and temporal modeling. Notably, our AnyCharV can preserve more detailed appearance and avoid lots of artifacts than Viggle and MIMO, especially looking at the generated arms and hands in Figure 3. Moreover, our approach can handle the complex human-object interactions very well, e.g., playing basketball, which can not be done with Make-a-Protagonist and Viggle. These results strongly affirm the effectiveness and robustness of our proposed AnyCharV.
Quantitative Results.
We show the results in Table 1. Our AnyCharV surpasses open-sourced Make-A-Protagonist by a significant margin, with 1.7% FVD, 51.2% DOVER++, and 2.49% CLIP Image Score improvements. Notably, AnyCharV performs better than MIMO by 8.08% regarding FVD, indicting the high reality and diversity of our method. Meanwhile, our proposed AnyCharV outperforms Viggle by 2.43% in terms of DOVER++ score, demonstrating the high aesthetic and technical quality of our generated videos. Considering Viggle and MIMO are both closed-source industrial products, our AnyCharV shows great effectiveness and robustness comparing with them. For inference cost, Make-A-Protagonist, Viggle, MIMO, and AnyCharV take 140, 2, 8, and 5 minutes to generate a 5 seconds 24 FPS video with resolution , respectively, further denoting the high efficiency of our approach.
User Study.
We conduct human evaluations to further evaluate our approach, comparing with three state-of-the-art methods. We gather 750 answers from 25 independent human raters, evaluating the identity preservation of the reference character, the motion consistency between the generated video and the target driving video, and the scene similarity and interaction. The average ranking is computed and shown in Table 2. It shows that our AnyCharV significantly outperforms the open-source Make-A-Protagonist and performs on par with the closed-source models Viggle and MIMO, demonstrating the effectiveness of our model.
4.3 Ablation Studies
We further do some ablation studies for our proposed framework, with deep analysis on self-boosting strategy, mask augmentation, and the choices of mask types.
Self-Boosting Strategy.
We introduce a self-boosting training mechanism to enhance the identity preservation. From Figure 4, we can learn that without self-boosting training, the identity and appearance of the reference character can not be well preserved, especially for the dressing cloth and face. Such performance drop can also be observed in Table 3, indicating the effectiveness of our self-boosting training.
Mask Augmentation.
AnyCharV augments the segmentation mask during the first stage to reduce the negative effect caused by the fine mask shape. The qualitative and quantitative results shown in Table 3 and Figure 4 demonstrate the necessity of our design. Without mask augmentation, the appearance of the reference character is disturbed greatly.
Mask Type.
We conduct ablation on mask types where we use different masks during stage 1 and stage 2. In the first training stage, the model composes the reference character and the target video scene within the accurate region, where the fine segmentation mask is preferred, indicated by the great spatial information loss for both character and background scene, produced by lines ‘Box & Seg.’ and ‘Box & Box’ in Figure 5 and Table 4. In the second training stage, the model is expected to better preserve the details of the reference character. In this case, we use more loose mask constrain, i.e., coarse bounding box mask, to guide the character generation, eliminating the adverse effect caused by the fine segmentation mask shape. Such performance improvement can be validated in Table 4 and Figure 5.
4.4 Application
Our AnyCharV can also be used to generate a video with a reference image generated by text-to-image (T2I) models, e.g., FLUX (Black Forest, 2023), driving by a target video generated by text-to-video (T2V) models, e.g., HunyuanVideo (Kong et al., 2024). As shown in Figure 6, we first utilize FLUX to generate a reference image and HunyuanVideo to synthesize a target video, then the generated target video is used to drive the synthesized reference character. These results clearly the strength and flexibility of our AnyCharV, proving its versatility.
5 Conclusion
In this work, we introduce a novel framework AnyCharV with fine-to-coarse guidance for controllable character video generation under a two-stage training strategy. The self-supervised composition strategy with fine mask guidance in the first stage basically learn to drive a reference image with a target driving video to guarantee the motion correctness and target scene maintenance. Self-boosting training is then carried out via building interactions between the reference and target characters with coarse mask guidance, where the detailed identity of the reference character can be better preserved. AnyCharV clearly beats state-of-the-art open-source models and performs just as well as leading closed-source industrial products. Most importantly, AnyCharV can be used for images and videos created by T2I and T2V models, showing its strong ability to generalize.
Appendix A Detailed Experimental Setting
A.1 Dataset
We build a character video dataset called CharVG to train our proposed model, which contains 5,482 videos from the internet with various characters. These videos range from 7 to 47 seconds in length. We utilize DWPose (Yang et al., 2023) to extract the whole-body poses from these videos, and the estimated poses are rendered as pose skeleton image sequences. Meanwhile, we use YOLOv8 (Reis et al., 2023) to obtain the segmentation and bounding box masks for the characters presented in these videos.
A.2 Implementation Details
The ReferenceNet and denoising UNet are both initialized from Stable Diffusion 1.5 (Rombach et al., 2022), where the motion module in the denoising UNet is initialized from AnimateDiff (Guo et al., 2024). The pose guider is initialized from ControlNet (Zhang et al., 2023a). In our training, firstly, we train the ReferenceNet, denoising UNet without motion module, and pose guider with individual frames from videos for 50,000 steps to learn the spatial information. Secondly, we perform self-boosted training on only denoising UNet without motion module with individual video frames for 3,000 steps to help the model eliminate the bad influence of the mask shape. Above two training processes are both conducted with resolution and batch size 64. Thirdly, we only train the motion module in the denoising UNet with 24-frame video clips with resolution for 10,000 steps to improve the temporal consistency. Last, the self-boosted training strategy is performed only on the denoising UNet with resolution for 10,000 steps to better preserve the identity of the reference character. Above two training processes are both conducted with batch size 8. All above are trained using learning rate 1e-5. During inference, the long video is generated sequentially with several short clips and temporally aggregated following (Tseng et al., 2023). We use a DDIM (Song et al., 2020) scheduler for 30 denoising steps with classifier-free guidance (Ho & Salimans, 2021) as 3.0. All of our experiments are finished on 8 NVIDIA H800 GPUs using PyTorch (Paszke et al., 2019). The overall training process takes about 3 days. Our model takes 5 minutes to generate a 5 seconds 24FPS video with resolution on a single NVIDIA H800 GPU.
A.3 Evaluation Metrics
We adopt several metrics to evaluate the generation quality. 1) FVD (Unterthiner et al., 2018) score is a widely used metric for assessing the generated videos. We compute FVD score between our generated videos and 1,000 real character videos from our dataset. 2) Dover++ (Wu et al., 2023a) score is a video quality assessment metric from both aesthetic and technical perspective, demonstrating the overall quality of the generated videos. 3) CLIP Image Score (Radford et al., 2021) is used to evaluate the similarity between the generated video and the reference character, validating the model capability for preserving the identity.
A.4 Comparison Methods
We compare our proposed AnyCharV with three state-of-the-art methods. We describe these methods as follows:
Make-A-protagonist (Zhao et al., 2023) memorizes the visual and motion information of the target video by fine-tuning the video generation model and guiding the generation of the desired output through masks and reference images during the inference stage.
Viggle (Viggle, 2024) performs rapid 3D modeling of the reference image and manipulates it according to the pose sequence of the input video, thereby accomplishing tasks such as character generation.
MIMO (Men et al., 2024) decomposes the character’s identity, pose, and scene to facilitate video synthesis.
Appendix B Additional Experimental Results
B.1 Qualitative Results
B.2 User Interface
We design a user interface to help human raters easily evaluate the generated videos. Our designed user interface is illustrated in Figure 9.
Appendix C Limitations and Future Works
While our AnyCharV model demonstrates impressive results in controllable character video generation, it may struggle when tasked with inferring the back view of a character from a front-facing reference image. Looking ahead, we plan to enhance the generalization capabilities of AnyCharV by integrating a more robust video generator and introducing more robust controls for open-world scenarios (Wang et al., 2024d; Che et al., 2024; Wang et al., 2024e).